OpenStudy (anonymous):

Find the slope of the line that passes through the points (1,4) and (4,7). a.-3/2 b. 1 c. 3/2 d. 2

jimthompson5910 (jim_thompson5910):

What are the y coordinates for the points given?

OpenStudy (anonymous):

this is the only information given to me on this problem

jimthompson5910 (jim_thompson5910):

If I gave you the point (2,3), what is the y coordinate of this point?

OpenStudy (anonymous):

oh sorry lol its 4 and 7....right?

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

Subtract the first from the second, so.. Second - First = 7 - 4 = 3

jimthompson5910 (jim_thompson5910):

Do the same (in the same order) for the x coordinates

OpenStudy (anonymous):

4-1=3

jimthompson5910 (jim_thompson5910):

so the slope is... Slope = (y difference)/(x difference) Slope = 3/3 Slope = 1

jimthompson5910 (jim_thompson5910):

It's coincidental in this case that the two differences are the same, usually they aren't

OpenStudy (anonymous):

okay thanks!:)

jimthompson5910 (jim_thompson5910):

sure thing